# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

### v3

@@ 5.1. Uniqueness Principle:
$D_{n}=\{x\in\mathbb{Z}^{d}:|x|_{1}\leq n\}$
$\partial D_{n}=\{x\in D:|x|_{1}=n\}$
and
$D_{n}^{\circ}=D\setminus\partial D_{n}.$
@@
  2. 2.
    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)+u(x)\quad\forall x\in D_{n}^{\circ}$.
@@
Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)+z(x)\quad\forall x\in D_{n}^{\circ}$ given by
$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)],$
where the expectation is taken assuming $X_{\mathscr{E}(\omega)}(0)=x\in D_{n}^{\circ}$. Furthermore, this function is also given alternately by
@@
for all $g:\mathcal{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem,
@@
Let $\mu$ be a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Then under assumptions 2.2 and 2.3, there exists a stationary ergodic measure $\lambda$ mutually absolutely continuous to $\mu$ under which $\{Y_{j}\}$ is stationary and ergodic.
To prove stationarity of the chain is enough to conclude ergodicity via arguments in [3]. Under 5.3,
$b=\int_{\mathcal{S}(d)}\mathscr{E(}\omega)d\lambda.$
@@ 5.2. Periodised Environments
Let $\mathbb{Z}^{d}/\sim=T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
@@
We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathcal{S}_{n}$ denote the set of all such $\omega$.
Fix $\omega\in\mathcal{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
@@
Denote by $R_{n}^{\omega}$ the resolvent given by
$R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega})}^{j}g(x)$
@@
As in the original paper by [7], in the following lemma we state a bound on the resolvent which in turn proves 5.3.
@@
There exists $c_{1}\equiv c_{1}(d)>0$ such that $\forall\omega\in\mathcal{S}_{n}$,
@@
To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [5]]
@@
Let the invariant measure for the random walk on the group $T_{n}$ given by $\{[X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}\}_{k\geq 0}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$. Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$. Therefore, one can check (or refer to Section 2 of [4]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
$\lambda_{n}(\tau_{x}\omega_{n})=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
Since $\mathcal{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
@@
for some measure $\lambda$ on $\mathcal{S}(d)$ and $K^{\mathscr{E}}_{d}$ is invariant under $\lambda$.
Now observe that
@@
$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\text{-a.e.}$
Since $\lambda_{n}$ extends as a measurable function over the entire $\mathcal{S}(d)$, we have that the family of functions
@@
$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty.$
Referring to arguments from [3], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
@@
Proof of Lemma 5.5 : Let $\omega\in\mathcal{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
@@
To see the equivalence, one notes the fact that $z(x)=Qf(x)$.
Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampère Operator for a given concave function $z$ given by
@@
$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x).$
Define the sets for $x\in D_{n}^{\circ}$
@@
$meas(I(x))\leq\frac{f(x)^{d}}{c(\mathscr{E},x)^{d}},$
where $meas(\cdot)$ is the Lebesgue measure on $\mathbb{R}^{d}$.
Now we state an easily provable lemma as in [7].
Lemma 6.3 (Lemma 8 of [7]).
@@
$\mathbb{E}_{x}\big[\sum_{j=0}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)]\leq\mathbb{E}_{x}\big[\sum_{j=0}^{\tau_{Kn}}g(X_{\mathscr{E}(\omega)}(j))\big]+\mathbb{E}_{x}\big[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)],$
where $x\in T_{n}$.
Using 6.1, one has that
@@
where $c^{\prime}\equiv c^{\prime}(d)$.
The last inequality follows due to the periodic nature of $g$.
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\Bigg]$ $=\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}>n^{2}\}}\Bigg]$ (2)
$\quad+\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].$
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
@@
$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
Since $|X_{\mathscr{E}(\omega)}(n)|_{1}^{2}$ is a submartingale, we have by Doob’s inequality that
$\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega)}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
@@
Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a natural proof of Theorem 3.5.
7. Proof of Theorem 5.1 and Theorem 3.5
@@
$\sum_{i=1}^{d}\mathscr{E}(\omega)(x_{0},e_{i})[z(x_{0}+e_{i})+z(x_{0}-e_{i})-2\beta]<-f(x_{0}).$
It is again easy to show that
@@
One also notes that for every $j\geq 1$,
$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
Letting $k\to\infty$ completes the proof.
Proof of Theorem 3.5: Fix any balanced elliptic environment $(\omega,\mathscr{E})$. Let $f:D_{n}\to[0,\infty)$ be a function. Observe that for every concave function $u\in\mathscr{A}(\mathscr{E},f)$ one has
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x).$
Denote by $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one, where it assigns 0.
Therefore, we have that
$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f).$
Denote by $\{X_{j}\}_{j\geq 0}$ the simple symmetric random walk. Then we have by Theorem 5.1 that
$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x).$
or by using the other form,
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]\geq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))].$
Similarly, we note that
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\geq\sum_{i=1}^{d}\Delta_{i}u(x).$
This shows that
$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right).$
As above, one can show that
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})].$
Therefore, we have established the following result.
Theorem 7.1.
Let $f:D_{n}\to[0,\infty)$ be a function. Let $(\omega,\mathscr{E})$ be a balanced elliptic environment. Then for all $x\in D_{n}^{\circ}$, we have the inequality,
$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}].$
One can now conclude Theorem 3.5 by taking the function $f:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$
$f(u)=\boldsymbol{1}_{x}(u)$
and letting $n\to\infty$.
8. Preparation for the proof of Theorem 3.7:
@@ 8.1. Action of Linear Transformations:
Fix $\omega\in\mathcal{S}(d)$. For this proof For a given linear transformation $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$, the action of $T$ on $\omega=\{\omega(x,v)\}_{x\in\mathbb{Z}^{d},v\in V_{d}}$ is given by
@@
$\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i}).$
@@
If $\tau_{X_{\omega}(n)}\omega$ has a stationary ergodic measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary and ergodic under $Q\circ T$.
@@
$\int_{\mathcal{S}(d)}K_{d}^{T}(\omega,B)dQ\circ T=\int_{\mathcal{S}(d)}K_{d}(T.\omega,T(B))dQ\circ T=Q\circ T(B)\quad\forall B\in\mathcal{B}({d})^{\prime}$
We shall use these notions fundamentally in our proof. We also note that showing stationarity of the local process is sufficient to exhibit its ergodicity, which we state in the following lemma. We shall prove the lemma by adapting arguments from [3].
@@
If $\mathbb{Q}$ is a measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary then it is also ergodic.
Let $\mathcal{S}^{\prime}=S(d)^{\mathbb{N}}$ denote all $S(d)$ valued trajectories of the local process. Let $T$ denote the canonical shifts or the $n^{th}$ iterate of the process. Denote by $\tilde{\mathcal{B}}$ the product $\sigma$-algebra. Let $A$ be a set which is invariant under $T$, i.e., $T^{-1}A=A$.
Let $h(\omega)=\tilde{P}_{\omega}(A)$ where $\tilde{P_{\omega}}$ is the law if the chain starts from $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}\sim\mathbb{Q}$. We must show
$\tilde{P}_{\mathbb{Q}}(A)=\int_{\mathcal{S}(d)}\tilde{P}_{\omega}(A)d\mathbb{Q}(\omega)=0\ \ or\ \ 1.$
Denote $\bar{\omega}(n)=\tau_{X_{\omega}(n)}\omega$.
Start by noting that under the canonical filtrations sequence $\mathscr{G}_{k}=\sigma(\omega,\bar{\omega}(1),\dots,\bar{\omega}(k-1))$,
$h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ T^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\text{-a.e.}$
@@
$h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\text{-a.e.}$ (3)
Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then $\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
@@
But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property and the fact that $\{h(\bar{\omega}(n)\}_{n\geq 0}$ is a martingale,
$\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\text{-a.e.}$ (4)
Combining 4 with ellipticity,
$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\text{-a.e.})\quad\forall\,e\in{V}_{d}.$ (5)
Iterating 5, we obtain
@@
$\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\text{\text{-a.e.}}$ (6)
@@
From 6 we get that $\mu(\tilde{B})=\mu(B)=0$ or $1$ (by ergodicity). From the mutual absolute continuity of $\mathbb{Q}$ and $\mu$, $\mathbb{Q}(B)=\tilde{P}_{\mathbb{Q}}(A)=0$ or $1$. This proves ergodicity of the local process in the environment $\mathbb{Q}$.

### v4

@@ 5.1. Uniqueness Principle:
$D_{n}$ $=\{x\in\mathbb{Z}^{d}:|x|_{1}\leq n\},$
$\partial D_{n}$ $=\{x\in D_{n}:|x|_{1}=n\},$
$D_{n}^{\circ}$ $=D_{n}\setminus\partial D_{n}.$
@@
  2. (2)
    $|Mz(x)|^{\frac{1}{d}}\geq\frac{f(x)}{c(\mathscr{E},x)}\quad x\in D_{n}^{\circ}$
@@
Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists a unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that
$|Mz(x)|^{\frac{1}{d}}=\frac{f(x)}{c(\mathscr{E},x)}.$
Furthermore, this function is also given alternately by
@@
for all $g:\mathscr{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem, which is a restatement of Theorem 3.1,
@@
Let $\mu$ be a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$. Then under assumptions 2.2 and 2.3, there exists a stationary ergodic measure $\lambda$ mutually absolutely continuous to $\mu$ under which $\{Y_{j}\}$ is stationary and ergodic.
To prove stationarity of the chain is enough to conclude ergodicity via arguments in [BolthausenSznitman2002]. Under 5.3,
$b=\int_{\mathscr{S}(d)}\mathscr{E(}\omega)d\lambda.$
@@ 5.2. Periodised Environments
Let $\mathbb{Z}^{d}/\sim\penalty 10000\ =T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
@@
We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathscr{S}_{n}$ denote the set of all such $\omega$.
Fix $\omega\in\mathscr{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
@@
For an environment $\omega$, denote by $R_{n}^{\omega}$ the resolvent given by
$R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega_{n}})}^{j}g(x)$
@@
As in the original paper by [Lawler1982], in the following lemma we state a bound on the resolvent which in turn proves 5.3.
@@
There exists $c_{1}\equiv c_{1}(d)>0$ such that $\forall\omega\in\mathscr{S}_{n}$,
@@
To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [Krengel1985]].
@@
Let the invariant measure for the random walk on the group $T_{n}$ given by $\{[X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}\}_{k\geq 0}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$. Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$. Therefore, one can check (or refer to Section 2 of [guo2012quenched]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
$\lambda_{n}=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
Since $\mathscr{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
@@
for some measure $\lambda$ on $\mathscr{S}(d)$ and $K^{\mathscr{E}}_{d}$ is invariant under $\lambda$. Now observe that
@@
[…]
for some constant $C$.
Since $\lambda_{n}$ extends as a measurable function over the entire $\mathscr{S}(d)$, we have that the family of functions
@@
$\int_{\mathscr{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<C^{\frac{p}{p-1}}.$
Referring to arguments from [BolthausenSznitman2002], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
@@
Proof of Lemma 5.5 : Let $\omega\in\mathscr{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
@@
Proof of Lemma 6.2: First we note the equivalence stated above. Recall the Monge-Ampère Operator for a given concave function $z$ given by
@@
$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x)=f(x)$
Therefore, one can show that for every $j\geq 1$,
$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]\leq\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
[…]
Define the sets for $x\in D_{n}^{\circ}$,
@@
$meas(I(x))=|Mz(x)|=\frac{f(x)^{d}}{c(\mathscr{E},x)^{d}},$
where $meas(\cdot)$ is the Lebesgue measure on $\mathbb{R}^{d}$. Now we state an easily provable lemma as in [Lawler1982].
Lemma 6.3 (Lemma 8 of [Lawler1982]).
@@
$\mathbb{E}_{x}\big[\sum_{j=0}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega_{n})}(j)\big)]\leq\mathbb{E}_{x}\big[\sum_{j=0}^{\tau_{Kn}}g(X_{\mathscr{E}(\omega_{n})}(j))\big]+\mathbb{E}_{x}\big[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega_{n})}(j)\big)],$
where $x\in T_{n}$. Using 6.1, one has that
@@
where $c^{\prime}\equiv c^{\prime}(d)$. The last inequality follows due to the periodic nature of $g$.
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\Bigg]$ $=\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}>n^{2}\}}\Bigg]$
$\quad+\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].$
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
@@
$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
Since $\{|X_{\mathscr{E}(\omega)}(k)|_{1}^{2}\}_{k\geq 1}$ is a submartingale, we have by Doob’s inequality that
$\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega_{n})}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
@@
Therefore, it is now enough to prove Theorem 5.1.
7. Proof of Theorem 5.1
@@
$\Big[\prod_{i=1}^{d}(2\beta-z(x+e_{i})-z(x-e_{i}))\Big]^{\frac{1}{d}}>\frac{f(x_{0})}{c(\mathscr{E},x_{0})}$
It is again easy to show that,
@@
8. Preparation for the proofs of Theorem 3.4:
@@ 8.1. Action of Linear Transformations:
Fix $\omega\in\mathscr{S}(d)$. For a given linear transformation $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$, the action of $T$ on $\omega=\{\omega(x,v)\}_{x\in\mathbb{Z}^{d},v\in V_{d}}$ is given by
@@
$\mathbb{P}_{0}^{\omega}(T(X_{\omega}(n))=T(X_{\omega}({n-1}))+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{\omega}(n)=X_{\omega}({n-1})+e_{i})=\omega(X_{n-1},e_{i}).$
@@
If $\tau_{X_{\omega}(n)}\omega$ has a stationary measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary under $Q\circ T$.
@@
$\int_{\mathscr{S}(d)}K_{d}^{T}(\omega,B)dQ\circ T=\int_{\mathscr{S}(d)}K_{d}(T.\omega,T(B))dQ\circ T=Q\circ T(B)\quad\forall B\in\mathcal{B}({d})^{\prime}$
We shall use these notions fundamentally in our proof. We also note that showing stationarity of the local process is sufficient to exhibit its ergodicity, which we state in the following lemma. We shall prove the lemma by arguments from [BolthausenSznitman2002].
@@
Let $\mathbb{Q}$ be a measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ is stationary. Further assume $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}=\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ in distribution under $\mathbb{Q}$. Then $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic under $\mathbb{Q}$.
Proof.
Stationarity follows trivially from the assumptions. We just have to show the ergodicity of the local process.
Let $\mathscr{S}^{\prime}=S(d)^{\mathbb{N}}$ denote all $S(d)$ valued trajectories of the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$.
Let $\theta$ denote the canonical shifts or the $n^{th}$ iterate of the process. Denote by $\tilde{\mathcal{B}}$ the product $\sigma$-algebra. Let $A$ be a set which is invariant under $\theta$, i.e., $\theta^{-1}A=A$.
Let $h(\omega)=\tilde{P}_{\omega}(A)$ where $\tilde{P}_{\omega}$ is the law if the chain starts from
$\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$ We must show
$\tilde{P}_{\mathbb{Q}}(A)=\int_{\mathscr{S}(d)}\tilde{P}_{\omega}(A)d\mathbb{Q}(\omega)\in\{0,1\}$
Let $P^{*}_{\omega}$ denote the law of the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ starting at $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$.
Therefore, it is enough to show,
${P}^{*}_{\mathbb{Q}}(A)=\int_{\mathscr{S}(d)}P^{*}_{\omega}(A)d\mathbb{Q}(\omega)\in\{0,1\}.$
Denote $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$.
Start by noting that under the canonical filtration sequence $\mathscr{G}_{k}=\sigma(\omega,\bar{\omega}(1),\dots,\bar{\omega}(k-1))$, $h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ\theta^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\text{-a.e.}$
@@
(2) $h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\text{-a.e.}$
Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then
$\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
@@
But this contradicts 2 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property, and the fact that $\{h(\bar{\omega}(n))\}_{n\geq 0}$ is a martingale,
(3) $\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}^{\Lambda}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\text{-a.e.}$
Combining 3 with ellipticity,
(4) $\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\text{-a.e.})\quad\forall\,e\in{V}_{d}.$
Iterating 4, we obtain
@@
(5) $\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\text{\text{-a.e.}}$
@@
From 5 we get that $\mu(\tilde{B})=\mu(B)=0$ or $1$ (by ergodicity). From the mutual absolute continuity of $\mathbb{Q}$ and $\mu$, $\mathbb{Q}(B)=P^{*}_{\mathbb{Q}}(A)=0$ or $1$. This proves ergodicity of the local process in the environment $\mathbb{Q}$. ∎